Imagine that you ask for a raise and your boss says that you can have one if you close at least half of your sales visits this month. You have 6 accounts and in the past the probability of closing one was 0.8. What is the probability that you get the raise?
Answer #1

Answer:

Given,

sample n = 6

p = 0.8

q = 1 - p

= 1 - 0.8

= 0.2

Consider binomial distribution

P(X = x) = nCx*p^x*q^(n-x)

P(X >= 3) = 1 - P(X < 3)

= 1 - [P(0) + P(1) + P(2)]

= 1 - [6C0*0.8^0*0.2^6 + 6C1*0.8^1*0.2^5 + 6C2*0.8^2*0.2^4]

= 1 - [0.000064 + 0.001536 + 0.01536]

= 1 - 0.01696

= 0.98304

Hence the probability that you get the raise is 0.98304